solve each system by substitution .5x-5y=19 and x-2y=-10
Answer by checkley71(8403)

X-2Y=-10 OR X=2Y-10 NOW SUBSTITUTING (2Y-10) FOR X IN THE OTHER EQUATION WE GET .5(2Y-10)-5Y=19 OR Y-5-5Y=19 OR -4Y=24 OR Y=24/-4 OR Y=-6
THE X-2(-6)=-10 OR X+12=-10 OR X=-10-12 OR X=-22
PROOF .5(-22)-5*-6=19 OR -11+30=19 OR 19=19